OpenStudy (anonymous):

will someone work this problem out with me solve by substitution method 4m+n=27 m-5n=54 im stuck and dont know what to do next

OpenStudy (anonymous):

m= 54+5n put in the first eq

OpenStudy (anonymous):

4(54+5n)+n=27

OpenStudy (anonymous):

21n=27-216=-189 n=-9 m=54+5(-9)=54-45=9

OpenStudy (anonymous):

i was taught n=27-4m m-5(27-4m)=54 m-135+20m=54 this is where i got stuck

OpenStudy (anonymous):

21m=54-135=-81 m = -81/21

OpenStudy (anonymous):

ahhh....we must get the same values of m and n....i guess i have done some mistake earliear:S

OpenStudy (anonymous):

oh yess got my mistake

OpenStudy (anonymous):

20n=-189

OpenStudy (anonymous):

can't you answer it with simultaneous equations?

OpenStudy (amistre64):

21m = 54+135...

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

off course we can do, but ann wanted by the substitution method

OpenStudy (amistre64):

missmys; there are alot of ways to solve it :)

OpenStudy (anonymous):

oh, ok :)

OpenStudy (amistre64):

the trick is trying to fit thru the right hoops at the right times for tinas teacher.... hes a tyrant ;)

OpenStudy (anonymous):

yes he is

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

ty everyone

OpenStudy (anonymous):

quick question why does the 20m change to 21m?

OpenStudy (amistre64):

20m + m = m(20+1) = 21m

OpenStudy (anonymous):

m+21m=21m

OpenStudy (anonymous):

coefficients are added up

OpenStudy (anonymous):

oh ok i missed that part

OpenStudy (amistre64):

if you ahve 20 apples on the counter and 1 apple on the table; how many apples you got all together? :)

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

uzma..... m+21m = 21m lol....youre funny :)

OpenStudy (amistre64):

so apparently i need a letter of recomendation to go to a 4 year college...

OpenStudy (amistre64):

who is sposed to writ eit?

OpenStudy (anonymous):

anyone that is in a professional state, like a guidence consular or someone like that

OpenStudy (anonymous):

UR MENTOR!!

OpenStudy (anonymous):

or form tutor

OpenStudy (amistre64):

i mentor myself lol

OpenStudy (anonymous):

??? i have a mentor at school but she only mentored me once pointless really :) what about form tutor then?

OpenStudy (amistre64):

whats a form tutor?

OpenStudy (amistre64):

at 40...there isnt much more that i can be mentored or tutored in ;)

OpenStudy (anonymous):

wait, do you go to a school? cause in our school we have our mentor and form tutor write those things

OpenStudy (anonymous):

form tutor just someone to talk to whenever you need it, but we have to go everyday and its pointless if you have no problems and you do discussions and stuff

OpenStudy (anonymous):

amistre64 is the tutor! or atleast for me when hes available he has never been wrong